OpenStudy (anonymous):

Given f(x) = –3x – 7, what is f(–4)? A. –1 B. –14 C. 5 D. 40

OpenStudy (anonymous):

\[f(-4) = -3(-4)-7\] You just plug in the -4 where x was originally.

OpenStudy (anonymous):

f(-4) = -3(-4)-7

OpenStudy (anonymous):

You just did what I did...what is -3(-4)-7 = ?

OpenStudy (anonymous):

\[-3 \times -4 -7 =?\]